Inspired by his recent trip to the Great​ Pyramids, Citibank trader Ruminder Dhillon wonders if he can make an intermarket arbitrage profit using Libyan dinars​ (LYD) and Saudi riyals​ (SAR). He has USD 1,000,000 to work with so he gathers the following quotes:

Citibank quotes U.S. dollar per Libyan dinar

$2.00 = LYD 1.00

National Bank of Kuwait quotes Saudi riyal per Libyan dinar

SAR 5.00 = LYD 1.00

Barclays quotes U.S. dollar per Saudi riyal

$0.25 = SAR 1.00

Is there an opportunity for an arbitrage​ profit? If yes, calculate the arbitrage profit.

Also clearly indicate the steps you would take to take advantage of the arbitrage opportunity.
Answer #1

Yes, arbitrage is possible

Fair rate between US Dollar and SAR = 2/5 = $0.4 per SAR

Since actual rate is not equal to fair rate, arbitrage is possible

Steps:

Convert USD 1,000,000 into SAR and get 1,000,000/0.25 = SAR 4,000,000

Convert into LYD = 4,000,000/5 = LYD 800,000

Convert into US Dollar = 800,000*2 = USD 1,600,000

Arbitrage profit = USD 1,600,000 – USD 1,000,000

= USD 600,000